the sum of john and harry's ages is 19 years. if the difference of their ages is 5 years,what are their ages?
Found 2 solutions by mananth, ewatrrr:
Answer by mananth(16946)

x+y=19
x-y=5
add the two equations
2x=19+5
2x=24
/2
x=12
y=7

Answer by ewatrrr(24785)

Hi,
Let x represent one of their age and y represent the other's age:
.
The question states the following to be true:
x -y = 5 years or x = y + 5
.
Sum of their ages is 19:
(y +5) + y = 19
.
Simplify and Solve:
2*y + 5 = 19
2*y = 14
y = 7 the other (x) is 12
.
Check Your answer:
7 + 12 = 19